Your brother starts out 10minutes before you on a bike path and rides at a rate of 9 miles per hour.If you ride at a rate of 12 miles per hour ,how long will it take you to catch up to him?
Answer by mananth(16949)

Your brother starts out 10minutes before you on a bike path and rides at a rate of 9 miles per hour.If you ride at a rate of 12 miles per hour ,how long will it take you to catch up to him?
..
brother speed = 9 mph
you speed = 12 mph
..
same direction one behind other.
so 12-9 = 3 mph is the catch up speed.
..
brther leaves 10 minutes early.
in 10 minutes he travels 10/60 * 9 = 3/2 miles
..
distance to catch up = 3/2 miles
catch up speed = 3 mph
catch up time = (3/2) / 3
= 1/2 hour
..
m.ananth@hotmail.ca